### v3

# Optimization of laser-driven proton acceleration in a near-critical-density plasma

Abstract
Optimizing laser and plasma parameters is crucial for enhancing accelerated proton energy in laser-driven proton acceleration with finite laser energy for applications such as cancer therapy. Tight focusing plays a significant role in improving laser-driven proton acceleration, which is generally believed as a result of the enhancement of laser intensity. However, we find that even at a fixed laser intensity, reducing the focal spot size still enhances the proton energy. Through particle-in-cell simulations and theoretical modeling, we find that at a small spot size (0.8 m), the maximum proton energy is enhanced by 56.3% compared to that obtained at a conventional spot size (3 m). This improvement is attributed to the dominance of ponderomotive-force-driven electrons at reduced spot sizes, which generate stronger charge-separation fields that propagate at higher velocities. Furthermore, to optimize proton acceleration, we analytically derive an ideal plasma density profile that promotes phase-stable proton acceleration, yielding an additional energy increase of 61.3% over the case of a tightly focused laser interacting with a planar target of uniform density. These findings remain robust under parameter variations, indicating that advanced focusing techniques combined with optimized plasma profiles could relax the demand for high laser energies, thereby reducing the reliance on large-scale laser facilities in medical and scientific applications.
I Introduction
Laser-driven plasma acceleration has emerged as a novel acceleration scheme in recent decades Mourou et al. (1998); Clark et al. (2000); Snavely et al. (2000); Daido et al. (2012), attracting extensive interdisciplinary interest due to its broad application prospects in both fundamental research and practical fields, including nuclear physics Ditmire et al. (1999); Roth et al. (2001), quantum electrodynamics Reichwein et al. (2025), materials science Boody et al. (1996); Conlon (1985), radiography Mackinnon et al. (2006); Cobble et al. (2002) and radiotherapy Tajima et al. (2009); Kroll et al. (2022); Li et al. (2023). Over the past 20 years, laser-driven proton acceleration has advanced rapidly, giving rise to a series of acceleration mechanisms such as Target Normal Sheath Acceleration (TNSA) Wilks et al. (2001); Pukhov (2001), Radiation Pressure Acceleration (RPA) Robinson et al. (2008); Yan et al. (2008); Yu et al. (2010), Collisionless Shock Acceleration (CSA) Silva et al. (2004); Dollar et al. (2012); Zhang et al. (2017) and Magnetic Vortex Acceleration (MVA) Bulanov et al. (2005, 2010); Bulanov (2022). With current laser technology, proton energy up to 150 MeV has been achieved from the interaction of intense lasers with solid targets Ziegler et al. (2024); Shou et al. (2025b), while Au ions have reached energies exceeding 1 GeV Wang et al. (2021); Martin et al. (2024). For advanced acceleration mechanisms, protons have been accelerated to over 90 MeV by a novel scheme associated with sheet crossing Shou et al. (2025a); Gong et al. (2020). Theoretically, GeV-level proton energies are predicted to be attainable at 1022 W/cm2 laser intensity Qiao et al. (2009); Chen et al. (2009); Bulanov et al. (2010). Moreover, laser driven wakefield proton acceleration represents a promising pathway for tens-of-GeV proton generationYi et al. (2013); Gong et al. (2024). Recently, laser-driven proton acceleration in near-critical density (NCD) plasma targets Seemann et al. (2024); Sharma (2018); Yogo et al. (2008) has gained attention as a promising approach for further energy enhancement. When an ultra-intense laser interacts with NCD plasma, relativistic motion increases the electron mass and reduces the plasma frequency. Once the plasma frequency drops below the laser frequency, the laser can penetrate the target, known as relativistic induced transparency (RIT) Jung et al. (2015); Powell et al. (2015); Palaniyappan et al. (2015). During this penetration, electrons experience enhanced heating via mechanisms such as direct laser acceleration (DLA) Pukhov et al. (1999); Liu et al. (2013), which finally improves proton energy.
Despite significant breakthroughs in acceleration mechanisms, current laser energies remain insufficient to accelerate protons beyond the 200 MeV threshold required for applications such as proton radiotherapyBulanov et al. (2008, 2002). Given that proton energy generally scales with laser energy Fuchs et al. (2006); Zimmer et al. (2021); Shen et al. (2021), a critical challenge lies in maximizing proton acceleration efficiency under present laser-energy constraints. Thus, optimizing laser and plasma parameters represents a critical avenue for enhancing proton energies, with laser focusing being one promising approach. Recent progress in wavefront correction has enabled near-diffraction-limited focusing, achieving a spot size of 1.1 m (full width at half maximum) in experiments Yoon et al. (2021). While tighter focusing is generally expected to boost proton energy through increased laser intensity, the influence of the focal‑spot size itself on energy enhancement has often been neglected.
In this paper, we demonstrate that reducing the laser focal spots of tightly focused lasers enhances proton energies in laser–near‑critical‑density plasma interactions, even at fixed laser intensity. This implies that higher proton energies can be achieved with lower laser energies. This enhancement originates from the longitudinal laser ponderomotive force, which scales inversely with the square of the laser spot size, becoming significantly stronger at a small focal spot, driving a more efficient electron acceleration. Consequently, a stronger and faster accelerating electric field is generated, leading to higher proton energies. Additionally, we design a down-ramp density profile to achieve velocity matching between protons and the accelerating electric field, further enhancing the proton energy. Through these optimizations, the proton energy can reach near-GeV level, meeting the threshold required for proton radiotherapy.
II Results
A series of two-dimensional (2D) particle-in-cell (PIC) simulations are conducted via relativistic fully self-consistent PIC code EPOCH Arber et al. (2015). A linearly polarized Gaussian laser pulse with a normalized amplitude of , a pulse duration of = 42 fs and a wavelength of = 800 nm is incident normally on a fully ionized hydrogen target with a density of 20 and a thickness of 7.5 m, where is the classical critical density. The simulation domain is a 40 m × 10 m rectangular box divided into 4000 × 1000 cells, with 10 macro-particles per cell per species. To validate our results, we also conduct three-dimensional (3D) simulations with a 40 m × 10 m × 10 m domain divided into 2000 × 500 × 500 cells and 4 macro-particles per cell per species. To investigate the effect of the laser spot size (defined by the laser field at the focus), we vary the laser spot size from 0.8 m to 5 m, with representative cases of 0.8 m and 3 m selected for detailed analysis.
II.1 Optimization by tightly focused laser
Figure 1(a) shows the temporal evolution of the maximum proton energy for 0.8 m and 3 m spot sizes, and the proton energy spectrum () at the end of the acceleration process is shown by the solid curve in Fig. 1(b). It can be observed that when the focal spot is 0.8 m, the cutoff energy of protons increases by 56.3% compared to the case with a 3 m spot, reaching 372 MeV. Although consuming a higher laser energy, the proton energy only reaches 238 MeV in the case with a larger focal spot. This counterintuitive result indicates that reducing the laser spot size, even at a fixed intensity and lower total laser energy, enhances proton acceleration significantly. The 3D results are depicted by the dashed lines in Fig. 1(b), confirming the validity of our findings. The proton cutoff energy in 3D simulations do not decrease much from 2D simulations, implying most of the proton energy is contributed by the Hole-Boring mechanism Robinson et al. (2009); Macchi et al. (2005); Palmer et al. (2011); Iwata et al. (2018). To elucidate the underlying physics, we analyze the electron density distributions at t = 24 for the 3 m case in Fig. 1(c) and the 0.8 m case in Fig. 1(d), where is the laser period. When the laser interacts with the NCD plasma, the ponderomotive force expels electrons forward and laterally, forming a channel without electrons and a density spike at the laser front, which can be seen in the axial electron density distribution near x = 0.8 m in Figs. 1 (c) and 1 (d). In the case with a focal point of 0.8 m in Fig. 1(d), a higher peak electron density can be observed, indicating a stronger laser ponderomotive pushing on electrons. The stronger ponderomotive force at the 0.8 m spot size results in a more intense electron acceleration, as further evidenced by the electron energy spectrum () represented by the dotted curves in Fig. 1(b).
To explain this observation, we propose a theoretical model. The transverse laser electric field follows a Gaussian distribution
where , is the spot size, is the Rayleigh length and r is the transverse distance. is the electric field at the focus at the 0.8 m spot and the 3 m spot. The longitudinal ponderomotive force at the axis
reaches its maximum value
at . The results clearly demonstrate that as the laser spot size decreases, the ponderomotive force, which scales inversely with the square of the spot size, becomes significantly stronger. This enhanced force drives a larger amount of electrons, consistent with the observation in Fig. 1(d) (0.8 m spot), where a more significant electron pile-up is observed compared to Fig. 1(c) (3 m spot). Additionally, for a large spot size, the ponderomotive force even can not reach its maximum value in the finite-thickness target.
To elucidate the proton acceleration process under different focal spot sizes, we present the accelerating electric field distributions along with the proton phase-space distributions at t = 40 and t = 60 in Figs. 2(a) and 2(b), respectively. At t = 40, the laser has not yet penetrated the target and the accumulated electrons driven by the laser ponderomotive force generate a charge-separation field, known as the Hole-Boring field. For the 0.8 m case, a smaller focal spot results in a stronger longitudinal ponderomotive force. Here, electrons are longitudinally pushed before being transversely expelled, leading to enhanced pile-up (Fig. 1(d)) and a larger charge-separation field, as shown in Fig. 2(a). Additionally, a stronger ponderomotive force induces a higher electron energy, causing the electron peak to propagate faster. This is evidenced by the greater displacement from the target front surface of the electron density peak in Fig. 1(d) (smaller spot) compared to Fig. 1(c) (larger spot) at t = 24. Consequently, the electric field in Fig. 2(a) exhibits faster propagation and higher amplitude for the smaller spot, efficiently improving the proton acceleration. At this stage, the TNSA field begins to form on the target rear side, as shown in Fig. 2(a). At t = 60, the Hole-Boring field has merged with the TNSA field, forming a distinct "peak-plateau" electric field structure (Fig. 2(b)). Protons pre-accelerated by the Hole-Boring field near the target front surface are then injected into the TNSA drifting field, gaining more energy through the subsequent drift acceleration. Protons at a small spot size are injected into the drifting sheath field earlier, before the sheath field decays or drifts away, leading to a higher proton energy.
To further clarify the proton energy disparity, Figs. 2(c) and 2(d) illustrate the proton acceleration process in the Hole-Boring and drifting field for 3 m and 0.8 m focal spots, respectively, with green lines marking the trajectories of the highest-energy protons. Obviously, the proton accelerated to the highest-energy is from a position more close to the target front surface for the 0.8 m focal spot in Fig. 2(d), indicating that high-energy protons are captured earlier by the accelerating field at a smaller spot, implying a long Hole-Boring acceleration. Comparing Fig. 2(c) and Fig. 2(d), it is obvious that the velocity of the Hole-Boring field at a small spot is larger than at a larger spot, confirming a stronger longitudinal ponderomotive force and faster electrons. As a result, these protons have gained higher energy before the laser penetrates the target. As they are injected into the drifting sheath field before its amplitude decays or it drifts away, protons at a small spot absorb more energy from the drifting field, achieving higher final energies. However, this enhancement occurs only when the spot size is sufficiently small (typically below 2 m). For focal spots larger than 3 m, where the effect of the ponderomotive force diminishes, proton acceleration reverts to being dominated by the available laser power. As a result, with increasing laser energy, electron heating becomes more intense. Consequently, at larger focal spots of over 3 m under the same laser intensity, the proton energy is generally higher when the focal spot is larger (Fig. 3).
To validate the robustness of these findings, we conducted parameter scans of spot sizes with varying target densities and laser intensities, as shown by scatters in Fig. 3. The results universally demonstrate enhanced proton energies with reduced spot sizes at a spot size smaller than 2 m. Furthermore, we propose a simple theoretical model to describe this phenomenon. The accelerating field is contributed by two distinct electron groups: electrons accelerated by the ponderomotive force at the laser front and electrons accelerated by the DLA mechanism in the channel. The field contributed by the ponderomotive-force accelerated electrons is in direct proportion to the ponderomotive force
where is the Lorentz factor and is a fit constant. The field contributed by electrons accelerated by the DLA mechanism is in direct proportion to the charge of electrons in the channel , where d is the target thickness. As the square velocity of the Hole-Boring field is , Macchi et al. (2005) the drifting field is assumed to have a square velocity . This assumption is from the continuity of accelerating field at the target rear surface in Fig. 2(d). As a result, the proton energy is . For simplicity, the effect of field velocity on proton energy is considered in the drifting field strength , which should be in direct proportion to . So,
where is a fit coefficient to describe the efficiency of accelerated channel electrons and is an area factor assuming the transverse radius of the field is 5 m. Then we have
Assuming an acceleration length of , the proton energy can be obtained as . This dual-component framework quantitatively explains the spot-size dependence of proton energy under various parameters. The theoretical predictions of our model are represented by solid lines in Fig. 3, which is consistent with simulation results (scatters) in the variation trend. As the model is rough, the coefficients are only set to fit the trend of spot size effects. is introduced to account for the collective effects of electrons and the role of ions, while compensates for the neglect of transverse effects. In Fig. 2(d), the highest-energy proton has exceeded the drifting field at the end of acceleration, so more optimization is in need to achieve the velocity matching between protons and field.
II.2 Optimization by down-ramp density target
Except for laser parameter optimization, proton energies can be further increased through target design, such as employing a density down-ramp Nakamura et al. (2010); Wan et al. (2019); Reichwein et al. (2021), a technique widely adopted in wakefield electron acceleration to optimize injection Hansson et al. (2015); Geddes et al. (2008); Martinez De La Ossa et al. (2017). The down-ramp density profile can be realized through plasma expansion induced by irradiating the rear surface of a target with a long weak laser pulse Tochitsky et al. (2020); Dulat et al. (2022). Due to the greater thickness of the target, the expansion of the rear surface is more significant than that of the front surface Dover et al. (2023). As a result, a target with a steep front surface and a down-ramp rear surface can be achieved. Alternatively, a stepped descending density profile can be formed by stacking multiple thin NCD targets with progressively decreasing densities. Under laser irradiation, such a stepped density distribution will be smoothed into a continuous down-ramp profile.
During the interaction of lasers with down-ramp near-critical-density targets, phase-stable proton acceleration can be achieved by matching the proton velocity to the accelerating electric field velocity. For linearly polarized lasers, the electric field velocity normalized by the light speed c in the Hole-Boring mechanism is given by Göthel et al. (2022)
where is the normalized laser amplitude and is the electron density. The amplitude of the electric field satisfies Göthel et al. (2022)
where and v is the velocity of the electric field as well as the proton velocity. Velocity matching requires the proton momentum to follow
where and . Solving this we can obtain
The electron density profile derived from this equation is plotted as the blue curve in Fig. 4(a). The left of Eq. 10 approximates , and then we get an approximate solution
as represented by the red curve in Fig. 4(a), where . The slope of this approximate solution is designed to be steeper than the numerically solved solution, as the laser plasma interaction may result in the expansion of the target and reshape the plasma profile.
To validate this velocity-matching acceleration, we conduct additional simulations to verify the enhancement of proton energy. For a fixed areal density and laser parameters, a 0.8 m focus laser is irradiated on a 1.5-m-thick target with a uniform electron density of 20 following this down-ramp. When a laser irradiates such a target, it first penetrates the uniform density region, exciting a strong electric field that imparts an initial velocity to the protons. Subsequently, the laser begins to interact with the down-ramp density region. As the laser propagates through the decreasing density profile, the electron density peak pushed by the laser radiation moves faster, thereby increasing the speed of the electric field to match the continuously increasing proton velocity. Protons are generated with a cutoff energy exceeding 600 MeV, a 61.3% increase over the interaction of a 0.8 m focus laser with a 7.5-m-thick target, as evidenced by the spectrum in Fig. 4(b).
Figures 4(c) and 4(d) illustrate the acceleration process with the proton velocity matching the accelerating field. Protons at the end of the accelerated beam experience stronger fields, continuously overtaking protons in front. This results in a distinct relatively-monoenergetic proton peak (about 50% energy spread) in Fig. 4(b), in contrast to the exponential spectrum in TNSA. As the field drifts, its velocity increases while the amplitude decays (Fig. 4(d)), maintaining the proton-field velocity matching throughout the acceleration progress. In detail, some high-energy protons exceed the field and leave the acceleration region, as shown in Fig. 4(d). The dotted line in this figure represents the front edge of the accelerating field, demonstrating high-energy protons exceeding the field. These protons are unable to experience the electric field and thus no longer gain energy until the electric field catches up with them again. Then, the field’s rising velocity allows it to catch up with these accelerated protons, enabling further proton acceleration.
III Conclusion
In summary, we demonstrate that reducing the laser spot size at fixed intensity enables higher proton energies despite a decrease in total laser energy. This unexpected enhancement is attributed to the electron energy enhancement at small spot sizes, where the ponderomotive force, which scales inversely with the square of the spot size, becomes significantly stronger. This stronger force produces a more intense and faster-propagating charge-separation field, leading to a more efficient proton acceleration. Meanwhile, we design a target with a density down-ramp to achieve velocity matching between protons and the accelerating field, enabling phase‑stable acceleration and a further increase in proton energy. Our results suggest that advances in laser‑focusing techniques can push proton energies to higher levels even with lower total laser energies, thereby reducing the reliance on large‑scale laser facilities. Together with optimized target down‑ramps, spot‑size minimization offers a practical pathway toward efficient laser‑plasma acceleration for applications that demand high particle energies from compact laser systems.
Acknowledgements.
This work was supported by the National Natural Science Foundation of China (Grant No. 11921006) and the National Grand Instrument Project (No. 2019YFF01014400). Deji Liu is supported by the China Postdoctoral Science Foundation (No. 2025M783355). The simulations are supported by the High-Performance Computing Platform of Peking University.References
- Contemporary particle-in-cell approach to laser-plasma modelling. Plasma Physics and Controlled Fusion 57 (11), pp. 1–26. Cited by: §II.
- Laser-driven ion source for reduced-cost implantation of metal ions for strong reduction of dry friction and increased durability. Laser and Particle Beams 14 (3), pp. 443–448. Cited by: §I.
- Advanced ion acceleration mechanisms. In 2022 IEEE Advanced Accelerator Concepts Workshop (AAC), pp. 1–6. Cited by: §I.
- Accelerating protons to therapeutic energies with ultraintense, ultraclean, and ultrashort laser pulses. Medical physics 35 (5), pp. 1770–1776. Cited by: §I.
- Generation of gev protons from 1 pw laser interaction with near critical density targets. Physics of plasmas 17 (4). Cited by: §I.
- Ion acceleration in a dipole vortex in a laser plasma corona. Plasma physics reports 31 (5), pp. 369–381. Cited by: §I.
- Oncological hadrontherapy with laser ion accelerators. Physics Letters A 299 (2-3), pp. 240–247. Cited by: §I.
- Enhanced collimated gev monoenergetic ion acceleration from a shaped foil target irradiated by a circularly polarized laser pulse. Phys. Rev. Lett. 103, pp. 024801. Cited by: §I.
- Measurements of energetic proton transport through magnetized plasma from intense laser interactions with solids. Physical Review Letters 84 (4), pp. 670. Cited by: §I.
- High resolution laser-driven proton radiography. Journal of applied physics 92 (4), pp. 1775–1779. Cited by: §I.
- Nuclear physics for materials technology accelerator based methods?. Contemporary physics 26 (6), pp. 521–558. Cited by: §I.
- Review of laser-driven ion sources and their applications. Reports on progress in physics 75 (5), pp. 056401. Cited by: §I.
- Nuclear fusion from explosions of femtosecond laser-heated deuterium clusters. Nature 398 (6727), pp. 489–492. Cited by: §I.
- Finite spot effects on radiation pressure acceleration from intense high-contrast laser interactions with thin targets. Physical review letters 108 (17), pp. 175005. Cited by: §I.
- Enhanced ion acceleration from transparency-driven foils demonstrated at two ultraintense laser facilities. Light: Science & Applications 12 (1), pp. 71. Cited by: §II.2.
- Subpicosecond pre-plasma dynamics of a high contrast, ultraintense laser–solid target interaction. Optics Letters 47 (21), pp. 5684–5687. Cited by: §II.2.
- Laser-driven proton scaling laws and new paths towards energy increase. Nature physics 2 (1), pp. 48–54. Cited by: §I.
- Plasma-density-gradient injection of low absolute-momentum-spread electron bunches. Physical review letters 100 (21), pp. 215004. Cited by: §II.2.
- Laser wakefield acceleration of ions with a transverse flying focus. Physical Review Letters 133 (26), pp. 265002. Cited by: §I.
- Proton sheet crossing in thin relativistic plasma irradiated by a femtosecond petawatt laser pulse. Physical Review E 102 (1), pp. 013207. Cited by: §I.
- Optimized laser ion acceleration at the relativistic critical density surface. Plasma Physics and Controlled Fusion 64 (4), pp. 044010. Cited by: §II.2, §II.2.
- Down-ramp injection and independently controlled acceleration of electrons in a tailored laser wakefield accelerator. Physical Review Special Topics—Accelerators and Beams 18 (7), pp. 071303. Cited by: §II.2.
- Plasma density limits for hole boring by intense laser pulses. Nature communications 9 (1), pp. 623. Cited by: §II.1.
- Scaling of ion energies in the relativistic-induced transparency regime. Laser and Particle Beams 33 (4), pp. 695–703. Cited by: §I.
- Tumour irradiation in mice with a laser-accelerated proton beam. Nature Physics 18 (3), pp. 316–322. Cited by: §I.
- Laser-driven time-limited light-sail acceleration of protons for tumor radiotherapy. Phys. Rev. Res. 5, pp. L012038. Cited by: §I.
- Generating overcritical dense relativistic electron beams via self-matching resonance acceleration. Physical review letters 110 (4), pp. 045002. Cited by: §I.
- Laser acceleration of ion bunches at the front surface of overdense plasmas. Physical review letters 94 (16), pp. 165003. Cited by: §II.1, §II.1.
- Proton radiography of a laser-driven implosion. Physical review letters 97 (4), pp. 045001. Cited by: §I.
- Narrow-band acceleration of gold ions to gev energies from ultra-thin foils. Communications Physics 7 (1), pp. 3. Cited by: §I.
- Optimizing density down-ramp injection for beam-driven plasma wakefield accelerators. Physical Review Accelerators and Beams 20 (9), pp. 091301. Cited by: §II.2.
- Ultrahigh-intensity lasers: physics of the extreme on a tabletop. Physics today 51 (1), pp. 22–28. Cited by: §I.
- High-energy ions from near-critical density plasmas via magnetic vortex acceleration. Physical review letters 105 (13), pp. 135002. Cited by: §II.2.
- Efficient quasi-monoenergetic ion beams from laser-driven relativistic plasmas. Nature communications 6 (1), pp. 10170. Cited by: §I.
- Monoenergetic proton beams accelerated by a radiation pressure driven shock. Physical review letters 106 (1), pp. 014801. Cited by: §II.1.
- Proton acceleration enhanced by a plasma jet in expanding foils undergoing relativistic transparency. New Journal of Physics 17 (10), pp. 103033. Cited by: §I.
- Particle acceleration in relativistic laser channels. Physics of Plasmas 6 (7), pp. 2847–2854. Cited by: §I.
- Three-dimensional simulations of ion acceleration from a foil irradiated by a short-pulse laser. Physical review letters 86 (16), pp. 3562. Cited by: §I.
- Stable gev ion-beam acceleration from thin foils by circularly polarized laser pulses. Phys. Rev. Lett. 102, pp. 145002. Cited by: §I.
- Plasma acceleration of polarized particle beams. Reports on Progress in Physics 88 (11), pp. 117001. External Links: Document, Link Cited by: §I.
- On the robustness of spin polarization for magnetic vortex accelerated proton bunches in density down-ramps. Plasma physics and controlled fusion 63 (8), pp. 085011. Cited by: §II.2.
- Relativistically correct hole-boring and ion acceleration by circularly polarized laser pulses. Plasma Physics and Controlled Fusion 51 (2), pp. 024004. Cited by: §II.1.
- Radiation pressure acceleration of thin foils with circularly polarized laser pulses. New journal of Physics 10 (1), pp. 013021. Cited by: §I.
- Fast ignition by intense laser-accelerated proton beams. Physical review letters 86 (3), pp. 436. Cited by: §I.
- Laser proton acceleration from a near-critical imploding gas target. Physical Review Letters 133 (2), pp. 025001. Cited by: §I.
- High energy electron and proton acceleration by circularly polarized laser pulse from near critical density hydrogen gas target. Scientific reports 8 (1), pp. 2191. Cited by: §I.
- Scaling laws for laser-driven ion acceleration from nanometer-scale ultrathin foils. Phys. Rev. E 104, pp. 025210. External Links: Document, Link Cited by: §I.
- Proton acceleration associated with sheet crossing in petawatt-laser-irradiated nanometer foils. Physical Review Letters 135 (21), pp. 215002. Cited by: §I.
- Laser-driven proton acceleration beyond 100 mev by radiation pressure and coulomb repulsion in a conduction-restricted plasma. Nature Communications 16 (1), pp. 1487. Cited by: §I.
- Proton shock acceleration in laser-plasma interactions. Physical Review Letters 92 (1), pp. 015002. Cited by: §I.
- Intense high-energy proton beams from petawatt-laser irradiation of solids. Physical review letters 85 (14), pp. 2945. Cited by: §I.
- Laser acceleration of ions for radiation therapy. Reviews of Accelerator Science and Technology 2 (01), pp. 201–228. Cited by: §I.
- Laser-driven collisionless shock acceleration of ions from near-critical plasmas. Physics of Plasmas 27 (8). Cited by: §II.2.
- Two-stage laser acceleration of high quality protons using a tailored density plasma. Physical Review Accelerators and Beams 22 (2), pp. 021301. Cited by: §II.2.
- Super-heavy ions acceleration driven by ultrashort laser pulses at ultrahigh intensity. Phys. Rev. X 11, pp. 021049. Cited by: §I.
- Energetic proton generation in ultra-intense laser–solid interactions. Physics of plasmas 8 (2), pp. 542–549. Cited by: §I.
- Generating high-current monoenergetic proton beams by a circularly polarized laser pulse in the phase-stable acceleration regime. Physical Review Letters 100 (13), pp. 135003. Cited by: §I.
- Proton acceleration by plasma wakefield driven by an intense proton beam. Laser and Particle Beams 31 (3), pp. 427–438. Cited by: §I.
- Laser ion acceleration via control of the near-critical density target. Physical Review E—Statistical, Nonlinear, and Soft Matter Physics 77 (1), pp. 016401. Cited by: §I.
- Realization of laser intensity over W/cm2. Optica 8 (5), pp. 630–635. Cited by: §I.
- Stable laser-driven proton beam acceleration from a two-ion-species ultrathin foil. Phys. Rev. Lett. 105, pp. 065002. Cited by: §I.
- Collisionless shock acceleration of high-flux quasimonoenergetic proton beams driven by circularly polarized laser pulses. Phys. Rev. Lett. 119, pp. 164801. Cited by: §I.
- Laser-driven high-energy proton beams from cascaded acceleration regimes. Nature Physics, pp. 1–6. Cited by: §I.
- Analysis of laser-proton acceleration experiments for development of empirical scaling laws. Physical Review E 104 (4), pp. 045210. Cited by: §I.